OpenStudy (anonymous):

Find the numerical value of the rational expression when x= 2 and x=-3 5x+6 / 2x^2+16x+24

OpenStudy (anonymous):

when x=2 answer is 1/4

OpenStudy (anonymous):

thank youuu!

OpenStudy (anonymous):

when x = 2, 5(2)+6/2(2)^2 + 16(2) + 24 = 16/64. which does simplify to 1/4. when x = -3, 5(-3)+6/2(-3)^2 + 16(-3) + 24 = -9/-6, which simplifies to 3/2.